Write down the equation of a pragressive wave of amplitude 0.02 m and period 0.05 s travelling along a streached string with a velocity 10 m/ sec

Grade:12

1 Answers

Khimraj
5 years ago
equation will y = Asin(wt - kx)
A = 0.02
T = 0.05
so w = 2\pi/T = 40\pi
v = 10
v = w/k 
10 = 40\pi/k
so k = 4\pi
then
y = 0.02sin(40\pit – 4\pix)
Hope it clears.